OpenStudy (anonymous):

State the degree: 5r^2s^3t^4 does that mean the first exponent (2) or the highest exponent (4)?

OpenStudy (anonymous):

It means the highest exponent.

OpenStudy (shane_b):

4th degree...since that's the highest exponent.

OpenStudy (anonymous):

Ok thanks :)

ganeshie8 (ganeshie8):

2+3+4

ganeshie8 (ganeshie8):

here r, s ,t are variables. you should take the sum of their powers. since its a a product

ganeshie8 (ganeshie8):

xy : degree = 2 x+y : degree = 1 x^2y + y : degree = 3 hope you get the idea

OpenStudy (shane_b):

@ganeshie8: Good catch...I didn't think of it being a product.

OpenStudy (shane_b):

For your first question, the degree of a polynomial expression is the degree of the term with the highest degree. For example, the degree of a term like x^2 is just 2. However, if the term is like x^2y, then that term has a degree of 3 because x^2y^1->2+1=3. Does that make more sense?

OpenStudy (anonymous):

oh I see yeah it makes sense! thank you! :)

OpenStudy (shane_b):

Cool :)